OpenStudy (anonymous):

Find the length of the curve given by r(t)=sqrt(2)/2t i+e^(t/2) j+e^(-t/2) z where -1<=t<=5.

OpenStudy (anonymous):

r'(t) = ( sqrt(2) / 2 , (1/2) e^(t/2) , (-1/2)e^(-t/2) )

OpenStudy (anonymous):

\[| r'(t)| = \sqrt{ (1/2) + (1/4)e^{t} +(1/4)e^{-t} }= \sqrt{ e^{-t} [ (1/2)e^{t} +(1/4)e^{2t} + (1/4) ]}\]

OpenStudy (anonymous):

\[|r'(t)|= \sqrt{ \frac{e^{-t}}{2} [ \frac{1}{2} (e^{t})^2 + e^{t} +\frac{1}{2}] } = \sqrt{ \frac{e^{-t}}{2} \times \frac{1}{2} (e^{t}+1)^2}\]

OpenStudy (anonymous):

\[|r'(t)| = \frac{1}{2} e^{- \frac{t}{2}} (e^{t} +1) = \frac{1}{2} ( e^{\frac{t}{2} } + e^{- \frac{t}{2}} )\]

OpenStudy (anonymous):

Therefore answer is simple

OpenStudy (anonymous):

\[= \frac{1}{2} \int\limits_{1}^{5} ( e^{\frac{t}{2}} + e^{- \frac{t}{2}} ) dt\]